OpenStudy (anonymous):

3/4 = -(x+11)/4x + 1/2x solve with work QUICK !

OpenStudy (anonymous):

PLEASEEE HELP !

OpenStudy (anonymous):

Multiply both sides by 4x to get: 3x = -(x + 11) + 4 3x = -x - 11 + 4 3x = -x - 7 Then, rearrange the terms.. 3x + x = -7 4x = -7 So x = -7/4

OpenStudy (anonymous):

thanks a ton

OpenStudy (anonymous):

a med